# FG-PE: Factor-graph Approach for Multi-robot Pursuit-Evasion

### v1

@@ IV-A4 Uncertainty Ellipse
Fig. 8 illustrates that pursuers reduce uncertainty by adjusting their movement. During the experiment and the planning of the pursuers, when a measurement is taken from the pursuers of the evader, the uncertainty decreases. While other approaches lack consideration of uncertainty in their predictions, the proposed method accounts for it and also minimizes it. In Fig. 9, the area of the uncertainty ellipse for different measurement frequencies is shown. As can be seen, decreasing the measurement frequency results in an increase in the area of uncertainty. The area is shown in log scale in Fig 9
Figure 9: The area of the uncertainty ellipse for different frequencies of measurements arriving from the pursuers. The
@@ IV-A5 Dropped messages
IV-B Hardware Real-World Simulation
To demonstrate that the provided solution is applicable in real-world scenarios, an experiment was designed to test it in practical situations. For this experiment, TurtleBots were utilized, and the development was conducted in ROS2. It shows the stability of the proposed solution even in real-world scenarios. Two pursuers and one evader were considered. The localization of the robots was based on the odometry of each robot, which increases the uncertainty and difficulty in localization and prediction.
@@ IV-B Hardware Real-World Simulation
As illustrated in Fig. 12, although the pursuers are far from the evader and the localization is uncertain, they are able to catch the evader. The speed of the pursuers in this experiment is twice the speed of the evader ($V_{p}=2V_{q}$).
V Conclusions
In this paper, we propose a factor graph-based approach to solve the pursuit-evasion (PE) problem, which is capable of both tracking and planning within dynamic environments. The pursuit-evasion problem is a critical challenge in robotics and multi-agent systems, where multiple pursuers must effectively track and capture an evader. Our proposed approach is robust against dropped messages, which is essential for real-time applications where communication may be unreliable. Additionally, it is designed to be scalable to accommodate different numbers of robots and obstacles, allowing it to be applicable in various operational contexts. To reduce optimization time, we utilize a single optimizer, which enhances scalability by reducing dimensionality and computational overhead. This design choice ensures that the algorithm remains efficient, even as the complexity of the environment increases. Furthermore, we incorporate additional factors into the factor graph to add more constraints, which enhances the accuracy of the estimation and tracking processes. This flexibility allows the system to adapt to changing conditions and maintain performance under various scenarios. The proposed method is rigorously compared with state-of-the-art techniques, demonstrating a significant reduction in both the time required to capture the evader and the average distance traveled by the pursuers. The conducted hardware experiments confirm the robustness of our approach, showcasing its effectiveness in real-world applications. Furthermore, scalability tests reveal that computation time decreases even as environmental complexity increases, indicating that our method can handle more demanding situations without sacrificing performance. In the future, we aim to integrate Gaussian processes into the solution to enable distributed message passing, which will further decrease the time needed to capture the evader. This integration could enhance the collaborative capabilities of the robots, allowing them to share information more efficiently and respond dynamically to changes in the environment. Another potential avenue for future work could involve solving the pursuit-evasion problem in the presence of dynamic obstacles and multiple evaders, which would significantly add complexity and realism to the pursuit-evasion scenario. By addressing these challenges, we aim to further improve the applicability of our approach in real-world situations where conditions are unpredictable and ever-changing.

### v2

@@ 4.1.4 Uncertainty Ellipse
Fig. 13 illustrates that pursuers reduce uncertainty by adjusting their movement. During the experiment and the planning of the pursuers, when a measurement is taken from the pursuers of the evader, the uncertainty decreases. While other approaches lack consideration of uncertainty in their predictions, the proposed method accounts for it and also minimizes it. In Fig. 14, the area of the uncertainty ellipse for different measurement frequencies is shown. As can be seen, decreasing the measurement frequency results in an increase in the area of uncertainty. The area is shown in log scale in Fig 14.
Figure 14: The area of the uncertainty ellipse for different frequencies of measurements arriving from the pursuers.
@@ 4.1.5 Dropped Messages
4.2 Real-world Experiment
To demonstrate that the provided solution is applicable in real-world scenarios, an experiment was designed to test it in practical situations. For this experiment, TurtleBots were utilized, and the development was conducted in ROS2. The experiment shows the performance of the proposed solution even in real-world scenarios. Two pursuers and one evader were considered. The localization of the robots was based on the odometry of each robot, which increases the uncertainty of localization and prediction.
@@ 4.2 Real-world Experiment
As illustrated in Fig. 17, although the pursuers are far from the evader. The speed of the pursuers in this experiment is twice the speed of the evader ($V_{p}=2V_{q}$). Videos of the experiments are available on the project website: https://sites.google.com/view/pursuit-evasion
4.3 Discussion
Utilizing factor graphs, our method (FG-PE) provides a unified framework that seamlessly combines estimation and planning, allowing pursuers to track the evader’s motion while choosing an efficient capture path. The advantages of FG-PE are rooted in the core properties of factor graphs and their robust architecture. The main reasons for the superior performance of FG-PE are:
1) FG-PE Explicitly Accounts for Uncertainty: The probabilistic nature of factor graphs allows FG-PE to directly model and minimize uncertainty in predictions. When a new measurement is introduced, the factor graph’s Jacobian matrix is updated to incorporate this richer information, leading to more accurate estimations and reduced uncertainty. As summarized in Sec. 4.1.4 and Fig. 14, increasing the frequency of measurement reduces the capture time by 16%. Moreover, the experiments show that for different measurement frequencies, the proposed solution can decrease the uncertainty area by a factor of $10^{5}$ on average.
2) FG-PE Handles Noise and Sparse Data: Factor graphs are highly effective at handling sparse and noisy measurements. This allows FG-PE to achieve accurate trajectory estimations and strategic planning even with fewer measurements, resulting in lower memory usage and computational overhead. Fig. 10 shows that, despite the presence of different odometry noise levels, which result in poorer estimates of the evader’s position, FG-PE is still able to generate effective plans that enable the pursuers to capture the evader.
3) FG-PE Provides a Flexible and Scalable Framework: The factor graph’s design allows it to adapt to varying numbers of pursuers and obstacles, making it easily scalable for multi-agent problems. As shown in Fig. 12, doubling the number of pursuers enables them to capture the evader faster, reducing the capture time by 55%.
4) FG-PE Maintains Robustness: The method is robust and performs well even when some communication messages are dropped, as the graph structure can manage these lost connections. As shown in Fig. 16, even when the fraction of dropped messages increases, the pursuers are still able to catch the evader. Doubling the fraction of dropped messages increases the capture time by 48%.
Overall, the ability of FG-PE to accurately estimate trajectories and plan strategically within a single and unified framework explains its consistent outperformance, especially in challenging speed scenarios.
5 Conclusions
In this paper, we propose a factor graph-based approach to solve the pursuit-evasion problem, which is capable of both tracking and planning within dynamic environments. The pursuit-evasion problem is a critical challenge in robotics and multi-agent systems, where multiple pursuers must effectively track and capture an evader. Our proposed approach is robust against dropped messages, which is essential for real-time applications where communication may be unreliable. Additionally, it is designed to be scalable to accommodate different numbers of robots and obstacles, allowing it to be applicable in various operational contexts.
To reduce optimization time, we utilize a single optimizer, which enhances scalability by reducing dimensionality and computational overhead. This design choice ensures that the algorithm remains efficient, even as the complexity of the environment increases. Furthermore, we incorporate additional factors into the factor graph to add more constraints, which enhances the accuracy of the estimation and tracking processes. This flexibility allows the system to adapt to changing conditions and maintain performance under various scenarios.
The proposed method is rigorously compared with state-of-the-art techniques, demonstrating a significant reduction in both the time required to capture the evader and the average distance traveled by the pursuers. The conducted hardware experiments confirm the robustness of our approach, showcasing its effectiveness in real-world applications. Furthermore, scalability tests reveal that computation time decreases even as environmental complexity increases, indicating that our method can handle more demanding situations without sacrificing performance.
In the future, we aim to integrate Gaussian processes into the solution to enable distributed message passing, which will further decrease the time needed to capture the evader. This integration could enhance the collaborative capabilities of the robots, allowing them to share information more efficiently and respond dynamically to changes in the environment. Another potential avenue for future work could involve solving the pursuit-evasion problem in the presence of dynamic obstacles and multiple evaders, which would significantly add complexity and realism to the pursuit-evasion scenario. By addressing these challenges, we aim to further improve the applicability of our approach in real-world situations where conditions are unpredictable and ever-changing.